# Multi-Step Reasoning with Large Language Models, a Survey

### v2

@@ 3. Step Generation, Evaluation and Control
Approach Domain (3.1) Step generation (3.2) Step evaluation (3.3) Step control Result Scratchpad (Nye et al., 2022) math word hand-wr/superv - greedy/1prompt PolyEval +19%, Python +21% Chain-of-thought (Wei et al., 2022b) math word hand-written - greedy/1prompt GSM8K +39%, SVAMP +10%, ASDiv +2%, AQuA +11%, MAWPS +14%, CSQA +1.8%, StrategyQA +0.2% ZS-CoT (Kojima et al., 2022) math word hand-written - greedy/1prompt MultiArith =89%, GSM8K =70% Auto-CoT (Zhang et al., 2023) math word model-generated - clustering MultiArith +0.3%, GSM8K +1%, AddSub +3.5%, AQuA +0.7%, SingleEq +0.4%, SVAMP +0.6%, CSQA +1%, StrategyQA +0%, Letter +0.7%, Coin +2.7% Complexity (Fu et al., 2022) math word hand-written self-consistency greedy/1prompt GSM8K +7%, MultiArith +3%, Penguins +3% Self-ask (Press et al., 2023) math word ext knowledge LLM multi-hop q. Bamboogle =60%, 2Wiki =40%, Musique =15% Self-verification (Weng et al., 2023) math word hand-written back-verify ensemble GSM8K +4%, SingleEq +2%, AddSub +4%, MultiArith +3%, AQuA +3%, SVAMP +1% Self-consistency (Wang et al., 2023b) math word hand-written majority ensemble GSM8K +18%, SVAMP +11%, AQuA +12%, StrategyQA +6%, ARC-c +4% Codex (Chen et al., 2021) code - tool-based - HumanEval =70% Self-debugging (Chen et al., 2024) code hand-written tool-based greedy Spider +9%, MBPP +12%, TransCoder +12% Fun-search (Romera-Paredes et al., 2024) code hand-written tool-based evolutionary alg cap set 8 =512 LLaMEa (van Stein and Bäck, 2024) code hand-written tool-based evolutionary alg BBOB +11% MathPrompter (Imani et al., 2023) math hand-written tool-based ensemble MultiArith =92% Program-of-thoughts (Chen et al., 2023) math word hand-wr, Codex Python+Consist. split reason/cmput GSM8K =71%, SVAMP =85%, ASDIV =85%, AddSub =92%, MultiArith = 99% Program-aided-lang (Gao et al., 2023) math word hand-wr, Codex NLP/Python ensemble GSM8K =72%, SVAMP =79%, ASDIV =79%, SingleEQ =96%, SingleOP =94%, AddSub = 92%, MultiArith = 99%, Penguins = 93% Refiner (Paul et al., 2024a) math word finetune critic model gen/crit feedback SVAMP =72%, GSM8K =78% Self-correction (Welleck et al., 2023) math word finetune corrector model gen/corr feedback MathProgSynth =24%, LexConstrGen =98%, ToxicityControl =0.0% Self-improvement (Huang et al., 2023) math word finetune self-assessment CoT/consistency GSM8K =82%, DROP =83%, ARC-c =90%, OpenBookQA =94%, ANLI-A3 =68% Say-can (Ahn et al., 2022) robot model-generated external model greedy Kitchen =31% Inner-monologue (Huang et al., 2022) robot hand-written various greedy TableTop =90%, Kitchen =60% Self-taught-reasoner (Zelikman et al., 2022) math word finetune augmentation greedy/feedback CommonsenseQA =72% Least-to-most (Zhou et al., 2023) math word hand-written self-assessment curriculum SCAN =99% Progressive-hint (Zheng et al., 2023) math word model-generated self-assessment stable prompt AddSub +2%, MultiArith +0%, SingleEQ +2%, SVAMP +3%, GSM8K +5%, AQuA +1% Self-refine (Madaan et al., 2023) math word model-generated self-assessment greedy/feedback Sentiment +32%, Dialog +49%, CodeOptim +8%, CodeRead +28%, MathReason +0%, AcronymGen +25%, ConstrainedGen +30% Tree-of-thoughts (Yao et al., 2024) puzzles model-generated self-assessment BFS/DFS Game24 =74%, CreativeWriting , Crossword Buffer-of-thoughts (Yang et al., 2024) math word thought template self-assessment buffer manager Game24 +11%, GeoShapes +20%, Checkmate +51% Algorithm-of-thoughts (Sel et al., 2023) puzzles model-generated self-assessment in-context RL GSM8K =89%, StrategyQA =84%, Crossword Beam-search (Xie et al., 2024) math word model-generated self-assessment Beam Search GSM8K +6%, AQuA +9%, StrategyQA +5% ReAct (Yao et al., 2023) action external knowledge self-assessment reinf learning ALFWorld =34%, WebShop =10% Reflexion (Shinn et al., 2024) decision model-generated ext model reinf learning HumanEval =91% Voyager (Wang et al., 2024a) Minecraft model-generated Minecraft reinf learning 15 x faster
@@
The second group, from Self-verification to Self-taught-reasoner, focuses on evaluation of the reasoning steps in the prompt. This column is shown in bold in the table. The approaches in this group aim to improve the Chain-of-thought results by reducing error accumulation that occurs when multiple steps are taken in a reasoning chain. A variety of step control methods is used by these approaches, which is discussed in more detail later. Note that not all approaches use natural language problems. For example, the subgroup of Codex to Program-aided-language focuses on formal languages. They generate code or math equations, typically in Python, to formalize the steps of the reasoning problem, or as result of the task. LLMs are quite good at code generation (Chen et al., 2021), and these approaches typically achieve good performance. The use of code also allows the approaches to call external programs such as interpreters and debuggers to evaluate the correctness of the reasoning steps that are generated.
@@ 3.1. Generation of Steps
If supervised learning can produce intermediate steps, would prompt learning be able to do so too?
@@ 3.1.2. Prompt using External Knowledge
Chain-of-thought prompts are written manually, by the researchers, an approach that does not scale. We can also use external information about the problem to improve the prompt. Press et al. (2023) study how subproblems are related to the main problem, which they call compositional reasoning. They study how often a model is able to answer the subproblems, but not the overall problem. This difference is called the compositionality gap. They find that in GPT-3, as model size increases, the single-hop question-answering performance improves faster than the multi-hop performance: while more powerful models memorize and recall more factual knowledge, no improvement in their compositional reasoning occurs. The ability to reason does not depend on the size of the model.
@@ 3.1.3. Model-Generated Prompt
Fu et al. (2022) introduce Complexity-based prompting. Inspired by Chain-of-thought and Self-consistency, their work specifically studies the impact of the complexity of the reasoning chain (the number of steps), and introduces a related reasoning approach (Complexity-based prompting). They find that prompts with the largest complexity perform best, and also that answers with the highest complexity are the best. Complexity-based prompting achieves somewhat higher performance on three math reasoning benchmarks: GSM8K improves 7 points, MathQA 6 points, and the Penguins benchmark from Big Bench Hard improve 3 percentage points.
@@ 3.2.1. Self-Assessment
A popular related approach is Self-consistency (Wang et al., 2023b). Self-consistency is a straightforward ensemble approach (a well-known machine learning technique to make a strong learner out of multiple weaker learners (Sagi and Rokach, 2018; Breiman, 2001)). Greedy single-path decoding is replaced by sampling diverse reasoning paths, evaluating them, and selecting the most consistent answer. Self-consistency asks the LLM to simply perform the same query multiple times, and takes the majority-vote of the answers, or decoding paths. Self-consistency works since complex reasoning problems typically allow different reasoning paths that lead to the correct answer. Figure 4 summarizes the approach.
Figure 4. Self-Consistency (Adapted from (Wang et al., 2023b)), showing majority voting
@@ 3.2.2. Tool-based Validation
Another approach to improve the accuracy of evaluating the reasoning steps is to switch from a natural to a formal language. The advantage of a formal language is that it is less ambiguous than a natural language. Examples are computer languages, such as Python, or mathematical equations. Using a formal language for reasoning is a popular approach, and we discuss seven papers. Many approaches generate the steps in Python, and the code can then be evaluated by a formal evaluator, such as a compiler, debugger, or interpreter.
@@
LLMs have been quite successful in generating computer code from natural language prompts. Chen et al. (2021) introduced Codex, a GPT model that was trained on publicly available code in the repository GitHub. A production version of this work was introduced under the name GitHub Copilot. Codex is able to generate correct programs from descriptions in natural language, such as commentary strings.
@@
The work on Codex is used as a basis for further research on reasoning in LLMs. Human programmers, when writing code, typically follow a cycle of writing some code, executing it to look for errors, and then using the feedback to improve the code. This step-by-step approach is followed in Self-debugging (Chen et al., 2024). It follows the same steps of code generation, code execution, and code explanation. Self-debugging is able to identify mistakes in its own code by investigating the execution results, and can also provide an explanation of the generated code, in natural language. It achieves strong performance: the text-to-SQL Spider benchmark improves 9 points, and the C++ to Python Transcoder benchmark improves by 12 percentage points.
@@
Several works generate working code tuned for solving specific problems automatically, without human feedback. Romera-Paredes et al. (2024) introduced FunSearch, an approach that integrates formal methods and LLMs to enhance mathematical reasoning and code generation. FunSearch uses a genetic approach with multiple populations of candidate solutions (programs), which are evaluated using a function depending on the problem specification. In addition to the evaluation function, also an initial program is given to the LLM in the first prompt. After evaluating a number of generated programs from the starting prompt, a new prompt is created, in an iterative fashion, combining a selection of sampled programs sorted according to their evaluation score, and the LLM is requested to generate a new program. Another work leverages evolutionary computation methods to generate and optimize evolutionary algorithms (van Stein and Bäck, 2024). This approach, LLaMEA (Large Language Model Evolutionary Algorithm), utilizes LLMs to design and optimize evolutionary algorithms. The approach uses LLMs to generate initial algorithmic structures, which are then refined through mutation and selection. This enhances the efficiency of algorithm design, particularly in fields requiring innovative and adaptive solutions, improving accuracy on the Black-Box Optimization Benchmark (Hansen and Ros, 2010) (BBOB) suite by 11 pecentage points. A key difference between FunSearch and LLaMEA is that LLaMEA uses a sample-efficient elitism strategy by iteratively improving the best-so-far solution, requiring significantly fewer prompt evaluations than the large-population strategy proposed in FunSearch. Evolutionary approaches let the LLM discover new algorithms, solving existing problems in new ways, or solving entirely new problems. Another method, Evolution-of-heuristics (Liu et al., 2024), was proposed for evolving code snippets for guided local search to solve combinatorial optimization problems, such as the Traveling Salesperson Problem.
@@
Two other approaches that use a formal language are Program-of-thought (Chen et al., 2023) and Program-aided-language (Gao et al., 2023). Both approaches use the LLM to generate Python and then use an interpreter to evaluate the result. The approaches are similar although Program-aided-language uses generic prompts, and has been tested on more benchmarks. Figure 3 (fourth column) illustrates the Program-aided-language approach. When the evaluation of the reasoning steps is offloaded to the Python interpreter, decomposing the natural language problem into executable code-steps remains the only task for the LLM. (Earlier work in math word problems showed how to decompose a problem and reach an answer (Ling et al., 2017).) Gao et al. (2023) provide extensive experimental evidence about the synergy between the neural LLM and the symbolic interpreter. Experiments are performed on 13 mathematical, symbolic, and algorithmic reasoning tasks, achieving more accurate results than much larger models (for example, Program-aided-language reported 72% on GSM8K and 93% on Penguins).
@@ 3.2.3. External Model Validation
We have seen many examples of succesful prompt-based in-context reasoning and evaluation (at inference time—where no parameters were changed). We will now look at reasoning approaches that follow a more traditional parameter training approach. All approaches evaluate the output of the model and generate corrective data. That data is then added to the training pipeline, and the model is subsequently finetuned.
@@
Welleck et al. (2023) follow a similar approach which they call Self-correction. Here, the corrector is a separate model specialized in refining the outputs of the generator. Unlike Refiner, where the generator is finetuned based on the critic, Self-correction finetunes the corrector to rectify errors in the hypotheses produced by the generator. Self-corrector is not applied to math word problems, but to program synthesis, where a small corrector reduces toxicity to 0%.
A third finetuning approach is Self-improvement (Huang et al., 2023). Here, too, the base model data is augmented by LLM-generated rationales, and then finetuned. They achieve 82% accuracy on GSM8K and similarly high scores on question answering and adversarial benchmarks. Noteworthy in all three finetuning approaches is that LLMs are capable of improving themselves by training on their own generated output, and that stability problems from feedback loops are overcome.

### v3

@@ 3. Step Generation, Evaluation and Control
Approach Domain (3.1) Step generation (3.2) Step evaluation (3.3) Step control Result Scratchpad (Nye et al., 2022) math word hand-wr/superv - greedy/1prompt PolyEval +19%, Python +21% Chain-of-thought (Wei et al., 2022b) math word hand-written - greedy/1prompt GSM8K +39%, SVAMP +10%, ASDiv +2%, AQuA +11%, MAWPS +14%, CSQA +1.8%, StrategyQA +0.2% ZS-CoT (Kojima et al., 2022) math word hand-written - greedy/1prompt MultiArith =89%, GSM8K =70% Auto-CoT (Zhang et al., 2023) math word model-generated - clustering MultiArith +0.3%, GSM8K +1%, AddSub +3.5%, AQuA +0.7%, SingleEq +0.4%, SVAMP +0.6%, CSQA +1%, StrategyQA +0%, Letter +0.7%, Coin +2.7% Complexity (Fu et al., 2022) math word hand-written self-consistency greedy/1prompt GSM8K +7%, MultiArith +3%, Penguins +3% Self-ask (Press et al., 2023) math word ext knowledge LLM multi-hop q. Bamboogle =60%, 2Wiki =40%, Musique =15% Self-verification (Weng et al., 2023) math word hand-written back-verify ensemble GSM8K +4%, SingleEq +2%, AddSub +4%, MultiArith +3%, AQuA +3%, SVAMP +1% Self-consistency (Wang et al., 2023c) math word hand-written majority ensemble GSM8K +18%, SVAMP +11%, AQuA +12%, StrategyQA +6%, ARC-c +4% Codex (Chen et al., 2021) code - tool-based - HumanEval =70% Self-debugging (Chen et al., 2024) code hand-written tool-based greedy Spider +9%, MBPP +12%, TransCoder +12% Fun-search (Romera-Paredes et al., 2024) code hand-written tool-based evolutionary alg cap set 8 =512 LLaMEa (van Stein and Bäck, 2024) code hand-written tool-based evolutionary alg BBOB +11% MathPrompter (Imani et al., 2023) math hand-written tool-based ensemble MultiArith =92% Program-of-thoughts (Chen et al., 2023) math word hand-wr, Codex Python+Consist. split reason/cmput GSM8K =71%, SVAMP =85%, ASDIV =85%, AddSub =92%, MultiArith = 99% Program-aided-lang (Gao et al., 2023) math word hand-wr, Codex NLP/Python ensemble GSM8K =72%, SVAMP =79%, ASDIV =79%, SingleEQ =96%, SingleOP =94%, AddSub = 92%, MultiArith = 99%, Penguins = 93% Refiner (Paul et al., 2024a) math word finetune critic model gen/crit feedback SVAMP =72%, GSM8K =78% Self-correction (Welleck et al., 2023) math word finetune corrector model gen/corr feedback MathProgSynth =24%, LexConstrGen =98%, ToxicityControl =0.0% Self-improvement (Huang et al., 2023) math word finetune self-assessment CoT/consistency GSM8K =82%, DROP =83%, ARC-c =90%, OpenBookQA =94%, ANLI-A3 =68% Say-can (Ahn et al., 2022) robot model-generated external model greedy Kitchen =31% Inner-monologue (Huang et al., 2022) robot hand-written various greedy TableTop =90%, Kitchen =60% Self-taught-reasoner (Zelikman et al., 2022) math word finetune augmentation greedy/feedback CommonsenseQA =72% Least-to-most (Zhou et al., 2023) math word hand-written self-assessment curriculum SCAN =99% Progressive-hint (Zheng et al., 2023) math word model-generated self-assessment stable prompt AddSub +2%, MultiArith +0%, SingleEQ +2%, SVAMP +3%, GSM8K +5%, AQuA +1% Self-refine (Madaan et al., 2023) math word model-generated self-assessment greedy/feedback Sentiment +32%, Dialog +49%, CodeOptim +8%, CodeRead +28%, MathReason +0%, AcronymGen +25%, ConstrainedGen +30% Tree-of-thoughts (Yao et al., 2024) puzzles model-generated self-assessment BFS/DFS Game24 =74%, CreativeWriting , Crossword Buffer-of-thoughts (Yang et al., 2024) math word thought template self-assessment buffer manager Game24 +11%, GeoShapes +20%, Checkmate +51% Algorithm-of-thoughts (Sel et al., 2023) puzzles model-generated self-assessment in-context RL GSM8K =89%, StrategyQA =84%, Crossword Beam-search (Xie et al., 2024) math word model-generated self-assessment Beam Search GSM8K +6%, AQuA +9%, StrategyQA +5% ReAct (Yao et al., 2023) action external knowledge self-assessment reinf learning ALFWorld =34%, WebShop =10% Reflexion (Shinn et al., 2024) decision model-generated ext model reinf learning HumanEval =91% Voyager (Wang et al., 2024a) Minecraft model-generated Minecraft reinf learning 15 x faster
@@
The second group, from Self-verification to Self-taught-reasoner, focuses on evaluation of the reasoning steps in the prompt. This column is shown in bold in the table. The approaches in this group aim to improve the Chain-of-thought results by reducing error accumulation that occurs when multiple steps are taken in a reasoning chain. A variety of step control methods is used by these approaches, which is discussed in more detail later. Note that not all approaches use natural language problems. For example, the subgroup of Codex to Program-aided-language focuses on formal languages. They generate code or math equations, typically in Python, to formalize the steps of the reasoning problem, or as result of the task. LLMs that are trained on computer code achieve good performance at code generation, Chen et al. (2021) report up to 77% accuracy on the HumanEval dataset. Gao et al. (2023) leverage this ability by translating problems into code, achieving between 61% and 99% accuracy on various mathematics tasks. The use of code also allows the approaches to call external programs such as interpreters and debuggers to evaluate the correctness of the reasoning steps that are generated.
@@ 3.1. Generation of Steps
If supervised learning can produce intermediate steps, would prompt learning be able to do so, too?
@@ 3.1.2. Prompt using External Knowledge
Chain-of-thought prompts are written manually, by the researchers, an approach that does not scale. We can use external information about the problem to improve the prompt. Press et al. (2023) study how subproblems are related to the main problem, which they call compositional reasoning. They study how often a model is able to answer the subproblems, but not the overall problem. This difference is called the compositionality gap. They find that in GPT-3, as model size increases, the single-hop question-answering performance improves faster than the multi-hop performance: while more powerful models memorize and recall more factual knowledge, no improvement in their compositional reasoning occurs. The ability to reason does not depend on the size of the model.
@@ 3.1.3. Model-Generated Prompt
Fu et al. (2022) introduce Complexity-based prompting. Inspired by Chain-of-thought and Self-consistency, their work specifically studies the impact of the complexity of the reasoning chain (the number of steps), and introduces a related reasoning approach (Complexity-based prompting). They find that prompts with the largest complexity perform best, and also that answers with the highest complexity are the best. Complexity-based prompting achieves somewhat higher performance on three math reasoning benchmarks: GSM8K improves 7 points, MathQA 6 points, and the Penguins benchmark from Big Bench Hard improves 3 percentage points.
@@ 3.2.1. Self-Assessment
A popular related approach is Self-consistency (Wang et al., 2023c). Self-consistency is a straightforward ensemble approach (a well-known machine learning technique to make a strong learner out of multiple weaker learners (Sagi and Rokach, 2018; Breiman, 2001)). Greedy single-path decoding is replaced by sampling diverse reasoning paths, evaluating them, and selecting the most consistent answer. Self-consistency asks the LLM to simply perform the same query multiple times, and takes the majority-vote of the answers, or decoding paths. Self-consistency works since complex reasoning problems typically allow different reasoning paths that lead to the correct answer. Figure 4 summarizes the approach.
Figure 4. Self-Consistency (Adapted from (Wang et al., 2023c)), showing majority voting over answers that the model produces
@@ 3.2.2. Tool-based Validation
Another approach to improve the accuracy of evaluating the reasoning steps is to switch from a natural to a formal language. The advantage of a formal language is that it is less ambiguous than a natural language. Examples are computer languages, such as Python, or mathematical equations. Using a formal language for reasoning is a popular approach, and we discuss seven papers. [^3] In addition to the languages discussed in this survey, there is a long tradition in AI of using formal reasoning based on logic. The interest in Chain-of-thought has also stimulated research into external symbolic logic solvers that are called by the LLM (Pan et al., 2023; Xu et al., 2024a; Saparov and He, 2023; Tafjord et al., 2020; Han et al., 2022; Srivastava et al., 2022; Zhong et al., 2022), or where LLMs transform the problem in a logic problem. The use of LLMs for generating formal languages is further discussed in section 4.4. Many approaches generate the steps in Python, and the code can then be evaluated by a formal evaluator, such as a compiler, debugger, or interpreter.
@@
LLMs have been quite successful in generating computer code from natural language prompts. Chen et al. (2021) introduced Codex, a GPT model that was trained on publicly available code in the repository GitHub. A production version of this work was introduced under the name GitHub Copilot. Codex is often able to generate correct programs from descriptions in natural language, such as commentary strings.
@@
The work on Codex is used as a basis for further research on reasoning in LLMs. Human programmers, when writing code, typically follow a cycle of writing some code, executing it to look for errors, and then using the feedback to improve the code. This step-by-step approach is followed in Self-debugging (Chen et al., 2024). It follows the same steps of code generation, code execution, and code explanation. Self-debugging is able to identify mistakes in its own code by investigating the execution results, and can also provide an explanation of the generated code, in natural language. It achieves strong performance: the text-to-SQL Spider benchmark improves by 9 points, and the C++ to Python Transcoder benchmark improves by 12 percentage points.
@@
Several works generate working code tuned for solving specific problems automatically, without human feedback. Romera-Paredes et al. (2024) introduced FunSearch, an approach that integrates formal methods and LLMs to enhance mathematical reasoning and code generation. FunSearch uses a genetic approach with multiple populations of candidate solutions (programs), which are evaluated using a function depending on the problem specification. In addition to the evaluation function, also an initial program is given to the LLM in the first prompt. After evaluating a number of generated programs from the starting prompt, a new prompt is created, in an iterative fashion, combining a selection of sampled programs sorted according to their evaluation score, and the LLM is requested to generate a new program. Another work leverages evolutionary computation methods to generate and optimize evolutionary algorithms (van Stein and Bäck, 2024). This approach, LLaMEA (Large Language Model Evolutionary Algorithm), utilizes LLMs to design and optimize evolutionary algorithms. The approach uses LLMs to generate initial algorithmic structures, which are then refined through mutation and selection. This enhances the efficiency of algorithm design, particularly in fields requiring innovative and adaptive solutions, improving accuracy on the Black-Box Optimization Benchmark suite (Hansen and Ros, 2010) (BBOB) by 11 percentage points. A key difference between FunSearch and LLaMEA is that LLaMEA uses a sample-efficient elitism strategy by iteratively improving the best-so-far solution, requiring significantly fewer prompt evaluations than the large-population strategy proposed in FunSearch. Evolutionary approaches let the LLM discover new algorithms, solving existing problems in new ways, or solving entirely new problems. Another method, Evolution-of-heuristics (Liu et al., 2024), was proposed for evolving code snippets for guided local search to solve combinatorial optimization problems, such as the Traveling Salesperson Problem.
@@
Two other approaches that use a formal language are Program-of-thought (Chen et al., 2023) and Program-aided-language (Gao et al., 2023). Both approaches use the LLM to generate Python and then use an interpreter to evaluate the result. The approaches are similar although Program-aided-language uses generic prompts, and has been tested on more benchmarks. Figure 3 (fourth column) illustrates the Program-aided-language approach. When the evaluation of the reasoning steps is off-loaded to the Python interpreter, decomposing the natural language problem into executable code steps remains the only task for the LLM. (Earlier work in math word problems showed how to decompose a problem and reach an answer (Ling et al., 2017).) Gao et al. (2023) provide extensive experimental evidence about the synergy between the neural LLM and the symbolic interpreter. Experiments are performed on 13 mathematical, symbolic, and algorithmic reasoning tasks, achieving more accurate results than much larger models (for example, Program-aided-language reported 72% on GSM8K and 93% on Penguins).
@@ 3.2.3. External Model Validation
We have seen many examples of successful prompt-based in-context reasoning and evaluation (at inference time—where no parameters were changed). We will now look at reasoning approaches that follow a more traditional parameter training approach. All approaches evaluate the output of the model and generate corrective data. That data is then added to the training pipeline, and the model is subsequently finetuned.
@@
Welleck et al. (2023) follow a similar approach, which they call Self-correction. Here, the corrector is a separate model specialized in refining the outputs of the generator. Unlike Refiner, where the generator is finetuned based on the critic, Self-correction finetunes the corrector to rectify errors in the hypotheses produced by the generator. Self-corrector is not applied to math word problems, but to program synthesis, where a small corrector reduces toxicity to 0%.
A third finetuning approach is Self-improvement (Huang et al., 2023). Here, too, the base model data is augmented by LLM-generated rationales, and then finetuned. The authors achieve 82% accuracy on GSM8K and similarly high scores on question answering and adversarial benchmarks. Noteworthy in all three finetuning approaches is that LLMs are capable of improving themselves by training on their own generated output, and that stability problems from feedback loops are overcome.